OpenStudy (anonymous):

. Janet Home went to Citizen Bank. She borrowed $7,000 at a rate of 8 percent. The date of the loan was September 20. Janet hoped to repay the loan on January 20. Assuming the loan is based on ordinary interest, Janet will pay back interest on January 20:

OpenStudy (jamesj):

So if she'd had the loan an entire year, the interest payment, I, would have been I = (% of interest)*(Principal) But she's had it only 3 months, so she should pay only 3/12 = 1/4 of that, as the interest did not at all compound. Hence in this case, I equals what?

OpenStudy (anonymous):

8 percent per year. Sep 20th to January 20th = 4 months. So, interest = 4/12 * 8 = 8/3 percent. At ordinary rate, interest amount = 7000 * 8/3 * 1/100 = ?

OpenStudy (jamesj):

yes, 4 months, my mistake.

OpenStudy (anonymous):

JamesJ - it is four months.

OpenStudy (anonymous):

i came up with 186.67?

OpenStudy (anonymous):

Sounds right.